OpenStudy (anonymous):

If 15m = 63, find the value of 4m. A. 16.8 B. 0.95 C. 4.2 D. 7.3

OpenStudy (anonymous):

16.8

OpenStudy (anonymous):

15m=63 m=63/15 4m=63(4)/15 4m=16.8

OpenStudy (anonymous):

thnkss